OpenStudy (brooklynriojas1):

find the probability of having 2,3, or 4 successes in 5 trials of a binomial experiment in which the probability of success is 40%. Round to the nearest tenth of a percent.

OpenStudy (holsteremission):

The probability of getting exactly \(k\) successes out of \(n\) trials with success probability \(p\) is given by \(\mathbb{P}(X=k)=\dbinom nk p^k(1-p)^{n-k}\). So the probability of getting \(2\), \(3\), or \(4\) successes (note that these are disjoint events) out of \(5\) trials and with probability of success \(p=\dfrac{2}{5}=40\%\) would be the summation of the probabilities of all of these events, or \[\mathbb{P}(X=2~|~X=3~|~X=4)=\sum_{k=2}^4\binom 5k\left(\frac{2}{5}\right)^k\left(\frac{3}{5}\right)^{5-k}\]